OpenStudy (anonymous):

Approximate the area under the curve of y = 1/x and above the x-axis from x = 1 to x = 2 by using a left sum with five subintervals of equal length. Will this be an under, or over-approximation to the actual area under the curve? Please explain.

OpenStudy (anonymous):

to answer the second question first, my drawing showing two subintervals shows that this will be an over-approximtion to the actual area under the curve

OpenStudy (anonymous):

ok got that part

OpenStudy (anonymous):

each subinterval will have a width of (2-1)/5=1/5 we have:\[A \approx f(1)\Delta x+ f(1.2)\Delta x + f(1.4) \Delta x + f(1.6) \Delta x + f(1.8) \Delta x\]where delta x=1/5

OpenStudy (anonymous):

\[A \approx (\frac{1}{1}+\frac{1}{1.2}+\frac{1}{1.4}+\frac{1}{1.6}+\frac{1}{1.8})\frac{1}{5}\]Do the computation and there you go :)

OpenStudy (anonymous):

got it.